OpenStudy (anonymous):

What is the name of the rule that allows this to be true? And is it actually true? If b/n = Log[a] then b = Log[a^n]

OpenStudy (anonymous):

And I guess also, what is the principle here that allows it to be true?

OpenStudy (anonymous):

Oh and does the base make any difference? If b/n = Log_E[a] then b = Log_E[a^n]

OpenStudy (joannablackwelder):

Yep, that is true. It is the exponent rule.

OpenStudy (joannablackwelder):



OpenStudy (joannablackwelder):

It is true for the same reason that when you raise an exponent to another exponent, they multiply.

OpenStudy (anonymous):

Got it, T/Y, So I could just quote 'the exponent rule' as an explanation for where this applies

OpenStudy (joannablackwelder):

I think the technical name is the logarithmic power rule.

OpenStudy (joannablackwelder):

yw :-)

OpenStudy (anonymous):

Ah I see, so Log power rule... if n=3 Log (a^n) = Log(a^3) = Log(a . a . a) = Log(a) + Log(a) + Log(a) = 3 Log(a) and because 1 = 3 Log(a) 1/3 = 3 Log(a) /3 1/3 = Log(a)

OpenStudy (joannablackwelder):

Yep :-)